# Topological and nontopological degeneracies in generalized string-net models

### v3

@@ VI.1 Rep($S_{3}$) and Vec($S_{3}$) categories
The fluxons are therefore $\mathbf{1},2,7$ and $8$, and, using Eq. (23), we can easily compute the degeneracies. Some examples are given in Table 5.
@@
[…]
@@
[…]
@@ VI.3 Haagerup category $\mathcal{H}_{3}$
The Haagerup category $\mathcal{H}_{3}$ is a good example of the universality of our formula: it is neither commutative, nor braided, nor Abelian, nor does it respect tetrahedral symmetry; see; e.g.; Ref. [24]. It has $N_{\mathcal{C}}=6$ simple objects $\{1,\alpha,\alpha^{*},\rho,\alpha\rho,\alpha^{*}\rho\}$ with quantum dimensions $\{1,1,1,d_{\rho},d_{\rho},d_{\rho}\}$, where $d_{\rho}=\frac{3+\sqrt{13}}{2}$. The Drinfeld center $\mathcal{H}_{3}$ contains $N_{\mathcal{Z}}=12$ simple objects $\{\mathbf{1},\mu^{1},\mu^{2},\mu^{3},\mu^{4},\mu^{5},\mu^{6},\pi_{1},\pi_{2},\sigma^{1},\sigma^{2},\sigma^{3}\}$ with quantum dimensions $\{1,3d_{\rho},3d_{\rho},3d_{\rho},3d_{\rho},3d_{\rho},3d_{\rho},3d_{\rho}+1,3d_{\rho}+2,3d_{\rho}+2,3d_{\rho}+2,3d_{\rho}+2\}$ so that the total quantum dimension is $\mathcal{D}=3(1+d_{\rho}^{2})$ [24]. The internal multiplicities are given by [50]
@@ A.1 Tubes $Q$ and their algebra $\mathcal{TA}$
The tubes can be composed (or multiplied): $Q^{i}_{rsj}Q^{k}_{s^{\prime}tl}$ means that $Q^{k}_{s^{\prime}tl}$ is stacked on top of $Q^{i}_{rsj}$ [see Fig. 11(b)]. The $\mathcal{TA}$ has the following structure [12]:
@@
and it is noncommutative in general. The coefficients in front of $Q^{m}_{rtn}$ in the sum on the right-hand side are the structure constants of the tube algebra. One can observe that the stacking is determined by the labels $r$ and $s$ of the open strings. Therefore, the tube algebra splits in different sectors labeled by $rs$. This will play a role later in the decomposition of the tube algebra.
Two important sets of tubes deserve particular attention:
@@
(i) The $N_{\mathcal{C}}$ tubes $Q_{11i}^{i}$ which have no open ends and are formed by “horizontal” closed strings [see Fig. 12(a)]. Each such tube corresponds to a simple object $i$ of $\mathcal{C}$. The $11$ sector consists only of these tubes ,and the tube algebra restricted to the $11$ sector reflects directly the algebra of the input category:
@@
(ii) The $N_{\mathcal{C}}$ tubes $Q^{1}_{rrr}$ formed by a “vertical” open string [see Fig. 12(b)], which correspond to the identity operator $\mathds{1}_{rr}$ in the $rr$ sector of $\mathcal{TA}$:
@@
Graphically, it is easy to see that stacking $Q^{1}_{rrr}$ on top of or below any tube of the same sector returns that second tube. The identity over $\mathcal{TA}$ is the sum over all of the $\mathds{1}_{rr}$:
@@ A.3 Simple idempotents and nilpotents $p$
and $\mathrm{Tr}\,p^{\alpha\alpha}=1$, meaning that the projected subspace is one-dimensional. The identity in the vector space is given by the sum over all these simple idempotents:
@@
it means that $|\alpha\rangle$ and $|\beta\rangle$ belong to the same irreducible block (say $A$) and that $p^{\alpha\beta}=|\alpha\rangle\langle\beta|=(p^{\beta\alpha})^{\dagger}$ (“off-diagonal elements”). Therefore, we denote as $p_{A}^{\alpha\alpha}$, $p_{A}^{\beta\beta}$, $p_{A}^{\alpha\beta}$, and $p_{A}^{\beta\alpha}$ the simple idempotents and nilpotents that belong to the same irreducible block.
@@ A.4 Minimal central idempotents $P_{A}$
The number of minimal central idempotents (topological sectors), $N_{\mathcal{Z}}$, is such that:
@@ A.5 Tube algebra sectors
If the input category $\mathcal{C}$ is noncommutative, the tube algebra in a given $rs$ sector need not be commutative. If it is noncommutative, it means that the corresponding vector space’s dimension $V^{rs}<N_{\mathcal{T}}^{rs}$ and that one needs to introduce an extra index $a$ to label orthonormal basis vectors. More generally, a simple idempotent is labeled as $p_{A}^{rr,aa}$ and a nilpotent as $p_{A}^{rs,ab}$ with $s\neq r$ or $b\neq a$ (meaning that $\alpha$ is actually $r,a$ with $r\in\mathcal{C}$ and $a=1,...,n_{A,r}$ here). We call $P_{A}^{rr}=\sum_{a}p_{A}^{rr,aa}$ the projector onto the quasiparticle $A$ in sector $rr$. (Our convention is that an idempotent denoted by a small $p$ is a simple idempotent that has a trace of $1$, whereas a capital $P$ is a projector with a trace that can be larger than $1$.) Also note that $P_{A}^{rr}$, in contrast to $P_{A}$, generally does not belong to the center of the tube algebra). It projects onto a subspace of dimension
@@ A.6 The $11$ sector and fluxons
In the $11$ sector, the number of tubes $N_{\mathcal{T}}^{11}=N_{\mathcal{C}}$ is larger than or equal to the vector space dimension $V^{11}$, which is larger than or equal to the number of fluxons $N_{\mathcal{F}}$:
@@ A.7 From tubes to idemnils: $M_{A}$
Finally, we express the tubes as linear combinations of the idemnils (this corresponds to writing the irreducible representations of the tubes in the basis spanned by the vectors $|A,r,a\rangle$),
@@ A.8 Fluxon degeneracies in the Levin-Wen model
The generalized Levin-Wen model has the property that the degeneracy of states is invariant under restructuring of the underlying graph by local (so-called “$F$”-) moves that preserve the numbers of vertices, edges, and plaquettes. In particular this means that to determine the number of states available by a plaquette, we can use the simplest plaquette possible, which is a single loop, connected to the rest of the graph by a single stem. The space of states available is then just the tube basis states $Q^{i}_{r1i}$ which has $i$ going around the loop and the stem labeled with $r$. We then decompose these basis states into idemnils via Eqs. (75) and (76). Since we are excluding any vertex violations, we consider values of $A$ which are fluxons only. For a particular value of $A$, the basis then consists of the orthonormal idemnils $p_{A}^{r1,ab}$ with $a=1,\ldots,n_{A,r}$ and $b=1,\ldots,n_{A,1}$. We can think of the factor $n_{A,1}$ as coming from the degeneracy associated with the “loop” end of the tube.
@@ A.9 Half-braidings $\Omega_{A}$
It is a simple matter to go from the projectors $P_{A}$’s to the string operators $W_{A}$’s or vice-versa. For example,
@@
Eventually, from $\Omega_{A}$ (or $M_{A}$), all properties of the Drinfeld center $\mathcal{Z}(\mathcal{C})$, such as the $S$, $T$, and fusion matrices, can also be easily obtained (see Eqs. (64), (65) and (60) in Ref. [12]).
@@ Appendix B Half-braidings for Rep($S_{3}$) and Vec($S_{3}$)
The Drinfeld center for Rep($S_{3}$) is given, e.g., in Ref. [49]. Some details about its tube algebra can be found in an Appendix in Ref. [53], and the half-braidings are derived in Ref. [54]. Since all $\Omega^{i}_{J,rsj}$ are one-dimensional, we write them under the form of a matrix $\Omega$ with $8$ rows indexed by particles $J=\{A,B,C,D,E,F,G,H\}$ and $17$ columns indexed by tubes $irsj=\{1111,2112,3113,1222,2221,3223,1333,2333,3331,3313,3133,3213,3123,3233,3323,3332,3333\}:$
@@
For Vec($S_{3}$), we have decomposed the tube algebra, and we give here the resulting half-braidings $\Omega_{J,rsj}^{i}$ as a matrix with $8$ rows indexed by $J$ as before and $36$ columns indexed by the tubes,
@@
As $n_{C,1}=2$ [see Eq. (38)], be aware that $\Omega_{C,11j}^{i}$ is actually a $2\times 2$ matrix of which Eq. (83) only gives the trace.
@@ Appendix C Fluxon identities
In Appendix A on the decomposition of the tube algebra, fluxons are identified as topological quasiparticles $A$ that have $n_{A,1}>0$. One can define a vector $\mathbf{n}_{1}$ with components $n_{A,1}$ with $A=1,...,N_{\mathcal{Z}}$ such that it has non-zero entries only when $A$ is a fluxon. Equations (11-12) are extra relations satisfied by the vector $\mathbf{n}_{1}$ that we call fluxon identities. Similar equations first appeared in the context of anyon condensation [42] and gapped boundaries [43, 44]. Equation (12) involving the $T$-matrix means that fluxons are necessarily bosons, i.e., they have a trivial twist $\theta_{A}=1$. When $\mathcal{Z}(\mathcal{C})$ is Abelian, Eq. (11) involving the $S$-matrix means that fluxons also have trivial mutual statistics [43]. When $\mathcal{Z}(\mathcal{C})$ is non-Abelian, the interpretation is less obvious. In the following, we provide a proof of the fluxon identites from the tube algebra, and we comment on the relation with anyon condensation.
@@ C.1.1 Commutative input category
In the commutative case, we can use the mock $S$-matrix $\tilde{s}$ to diagonalize simultaneously all the fusion matrices $N_{i}$. It is a unitary matrix but, unlike the $S$-matrix, it is not symmetric in general (a special case is when the input category is modular, in which case $\tilde{s}$ is a genuine $S$-matrix). Naturally, one would label the matrix elements as $\tilde{s}_{i,j}$ with $i$ and $j\in\mathcal{C}$. However, $\tilde{s}$ is also a unitary transformation between input strings and fluxons (see Eqs. (85) and (86) below). As such, physically it makes more sense to write a matrix element of $\tilde{s}$ as $\tilde{s}_{i,A}$ with rows indexed by input labels $i=1,..,N_{\mathcal{C}}$ and columns by fluxons $A\in\mathcal{F}$ with $N_{\mathcal{F}}=N_{\mathcal{C}}$. The matrix $\tilde{s}$ can be chosen such that its first row only contains strictly positive elements, in which case, $\tilde{s}_{1,A}=\sqrt{d_{A}/\mathcal{D}}=\sqrt{S_{\mathbf{1},A}}$. The mock $S$-matrix is not yet unique as its columns can always be permuted. By convention, we choose that the first column corresponds to the vacuum $A=\mathbf{1}$ of the output category so that $\tilde{s}_{i,\mathbf{1}}=d_{i}/\sqrt{\mathcal{D}}$. Examples of mock $S$-matrix are given in Secs. VI.1, VI.2 and VI.4.
@@
The action of the modular matrices $S$ and $T$ on the tubes is well-known: they act trivially on the empty tube (see Sec. 28.3 in [22]) so that
@@ C.1.2 Noncommutative input category
In the noncommutative case, we also have that the tube algebra restricted to the $11$ sector is the fusion algebra of the input category and that it decouples from the rest of the tube algebra. However, it is noncommutative and there is no mock-$S$ matrix. A consequence is that the number of tubes $N_{\mathcal{C}}$ is strictly larger than the vector space dimension $V_{11}$, which is strictly larger than the number of fluxons $N_{\mathcal{F}}$.
@@
as $(M^{1}_{A,111})_{a,b}=(\Omega^{1}_{A,111})_{a,b}=\delta_{a,b}$ (see above Eq. (40) in Ref. [14] or below Eqs. (52) and (53) in Ref. [12]). Therefore $Q_{1}$ is also the projector onto the $11$ sector, i.e. the empty tube $Q^{1}_{111}$, from which the fluxon identities (92) follow as in the previous subsection.

### v4

@@ VI.2 Tambara-Yamagami category TY3
The fluxons are therefore $\mathbf{1},2,7$ and $8$ and, using Eq. (23), we can easily compute the degeneracies. Some examples are given in Table 5.
@@
Table 5: Degeneracies of the $q^{th}$ excited state of a string net built from TY3 for various surface topologies up to $q=4$. Here $g$ is the genus and $b$ is the number of boundaries.
@@
The fusion matrices of TY3 are simultaneously diagonalized by the following mock $S$-matrix:
@@ VI.3 Haagerup category $\mathcal{H}_{3}$
The Haagerup category $\mathcal{H}_{3}$ is a good example of the universality of our formula: it is neither commutative, nor braided, nor Abelian, nor does it respect tetrahedral symmetry, see e.g. [24]. It has $N_{\mathcal{C}}=6$ simple objects $\{1,\alpha,\alpha^{*},\rho,\alpha\rho,\alpha^{*}\rho\}$ with quantum dimensions $\{1,1,1,d_{\rho},d_{\rho},d_{\rho}\}$ where $d_{\rho}=\frac{3+\sqrt{13}}{2}$. The Drinfeld center $\mathcal{H}_{3}$ contains $N_{\mathcal{Z}}=12$ simple objects $\{\mathbf{1},\mu^{1},\mu^{2},\mu^{3},\mu^{4},\mu^{5},\mu^{6},\pi_{1},\pi_{2},\sigma^{1},\sigma^{2},\sigma^{3}\}$ with quantum dimensions $\{1,3d_{\rho},3d_{\rho},3d_{\rho},3d_{\rho},3d_{\rho},3d_{\rho},3d_{\rho}+1,3d_{\rho}+2,3d_{\rho}+2,3d_{\rho}+2,3d_{\rho}+2\}$ so that the total quantum dimension is $\mathcal{D}=3(1+d_{\rho}^{2})$ [24]. The internal multiplicities are given by [50]
@@ A.1 Tubes $Q$ and their algebra $\mathcal{TA}$
The tubes can be composed (or multiplied): $Q^{i}_{rsj}Q^{k}_{s^{\prime}tl}$ means that $Q^{k}_{s^{\prime}tl}$ is stacked on top of $Q^{i}_{rsj}$ [see Fig. 11(b)]. The $\mathcal{TA}$ has the following structure [12]
@@
and is noncommutative in general. The coefficients in front of $Q^{m}_{rtn}$ in the sum on the right-hand side are the structure constants of the tube algebra. One can observe that the stacking is determined by the labels $r$ and $s$ of the open strings. Therefore, the tube algebra splits in different sectors labeled by $rs$. This will play a role later in the decomposition of the tube algebra.
Two important sets of tubes deserve a particular attention:
@@
(1) the $N_{\mathcal{C}}$ tubes $Q_{11i}^{i}$ which have no open ends and are formed by “horizontal” closed strings [see Fig. 12(a)]. Each such tube corresponds to a simple object $i$ of $\mathcal{C}$. The $11$ sector consists only of these tubes and the tube algebra restricted to the $11$ sector reflects directly the algebra of the input category:
@@
(2) The $N_{\mathcal{C}}$ tubes $Q^{1}_{rrr}$ formed by a “vertical” open string [see Fig. 12(b)] which correspond to the identity operator $\mathds{1}_{rr}$ in the $rr$ sector of $\mathcal{TA}$:
@@
Graphically, it is easy to see that stacking $Q^{1}_{rrr}$ on top or below of any tube of the same sector returns that second tube. The identity over $\mathcal{TA}$ is the sum over all of the $\mathds{1}_{rr}$:
@@ A.3 Simple idempotents and nilpotents $p$
and $\mathrm{Tr}\,p^{\alpha\alpha}=1$, meaning that the projected subspace is one dimensional. The identity in the vector space is given by the sum over all these simple idempotents:
@@
it means that $|\alpha\rangle$ and $|\beta\rangle$ belong to the same irreducible block (say $A$) and that $p^{\alpha\beta}=|\alpha\rangle\langle\beta|=(p^{\beta\alpha})^{\dagger}$ (“off-diagonal elements”). Therefore, we denote $p_{A}^{\alpha\alpha}$, $p_{A}^{\beta\beta}$, $p_{A}^{\alpha\beta}$ and $p_{A}^{\beta\alpha}$ the simple idempotents and nilpotents that belong to the same irreducible block.
@@ A.4 Minimal central idempotents $P_{A}$
The number of minimal central idempotents (topological sectors), $N_{\mathcal{Z}}$ is such that:
@@ A.5 Tube algebra sectors
If the input category $\mathcal{C}$ is noncommutative, the tube algebra in a given $rs$ sector need not be commutative. If it is noncommutative, it means that the corresponding vector space’s dimension $V^{rs}<N_{\mathcal{T}}^{rs}$ and that one needs to introduce an extra index $a$ to label orthonormal basis vectors. More generally, a simple idempotent is labeled as $p_{A}^{rr,aa}$ and a nilpotent as $p_{A}^{rs,ab}$ with $s\neq r$ or $b\neq a$ (meaning that $\alpha$ is actually $r,a$ with $r\in\mathcal{C}$ and $a=1,...,n_{A,r}$ here). We call $P_{A}^{rr}=\sum_{a}p_{A}^{rr,aa}$ the projector onto the quasiparticle $A$ in sector $rr$ (our convention is that an idempotent denoted by a small $p$ is a simple idempotent that has a trace of $1$, whereas a capital $P$ is a projector with a trace that can be larger than $1$. Also note that $P_{A}^{rr}$, in contrast to $P_{A}$, does generally not belong to the center of the tube algebra). It projects onto a subspace of dimension
@@ A.6 The $11$ sector and fluxons
In the $11$ sector, the number of tubes $N_{\mathcal{T}}^{11}=N_{\mathcal{C}}$ is larger or equal than the vector space dimension $V^{11}$ which is larger or equal than the number of fluxons $N_{\mathcal{F}}$:
@@ A.7 From tubes to idemnils: $M_{A}$
Finally, we express the tubes as linear combinations of the idemnils (this corresponds to writing the irreducible representations of the tubes in the basis spanned by the vectors $|A,r,a\rangle$) :
@@ A.8 Fluxon degeneracies in the Levin-Wen model
The generalized Levin-Wen model has the property that the degeneracy of states is invariant under restructuring of the underlying graph by local (so-called “$F$”-) moves that preserve the numbers of vertices, edges, and plaquettes. In particular this means that to determine the number of states available by a plaquette, we can use the simplest plaquette possible, which is a single loop, connected to the rest of the graph by a single stem. The space of states available is then just the tube basis states $Q^{i}_{r1i}$ which has $i$ going around the loop and the stem labeled with $r$. We then decompose these basis states into idemnils via Eqs. (75) and (76). Since we are excluding any vertex violations we consider values of $A$ which are fluxons only. For a particular value of $A$, the basis then consists of the orthonormal idemnils $p_{A}^{r1,ab}$ with $a=1,\ldots,n_{A,r}$ and $b=1,\ldots,n_{A,1}$. We can think of the factor $n_{A,1}$ as coming from the degeneracy associated with the “loop” end of the tube.
@@ A.9 Half-braidings $\Omega_{A}$
It is a simple matter to go from the projectors $P_{A}$’s to the string operators $W_{A}$’s or vice-versa. For example
@@
Eventually, from $\Omega_{A}$ (or $M_{A}$), all properties of the Drinfeld center $\mathcal{Z}(\mathcal{C})$, such as the $S$, $T$ and fusion matrices, can also be easily obtained (see Eqs. (64), (65) and (60) in Ref. [12]).
@@ Appendix B Half-braidings for Rep($S_{3}$) and Vec($S_{3}$)
The Drinfeld center for Rep($S_{3}$) is given, e.g., in Ref. [49]. Some details about its tube algebra can be found in an Appendix in Ref. [53] and the half-braidings are derived in Ref. [54]. Since all $\Omega^{i}_{J,rsj}$ are one-dimensional, we write them under the form of a matrix $\Omega$ with $8$ rows indexed by particles $J=\{A,B,C,D,E,F,G,H\}$ and $17$ columns indexed by tubes $irsj=\{1111,2112,3113,1222,2221,3223,1333,2333,3331,3313,3133,3213,3123,3233,3323,3332,3333\}:$
@@
For Vec($S_{3}$), we have decomposed the tube algebra and give here the resulting half-braidings $\Omega_{J,rsj}^{i}$ as a matrix with $8$ rows indexed by $J$ as before and $36$ columns indexed by the tubes
@@
As $n_{C,1}=2$ [see Eq. (38)], beware that $\Omega_{C,11j}^{i}$ is actually a $2\times 2$ matrix of which Eq. (83) only gives the trace.
@@ Appendix C Fluxon identities
In Appendix A on the decomposition of the tube algebra, fluxons are identified as topological quasiparticles $A$ that have $n_{A,1}>0$. One can define a vector $\mathbf{n}_{1}$ with components $n_{A,1}$ with $A=1,...,N_{\mathcal{Z}}$ such that it has non-zero entries only when $A$ is a fluxon. Equations (11-12) are extra relations satisfied by the vector $\mathbf{n}_{1}$ that we call fluxon identities. Similar equations first appeared in the context of anyon condensation [42] and gapped boundaries [43, 44]. Equation (12) involving the $T$-matrix means that fluxons are necessarily bosons, i.e., they have a trivial twist $\theta_{A}=1$. When $\mathcal{Z}(\mathcal{C})$ is Abelian, Eq. (11) involving the $S$-matrix means that fluxons also have trivial mutual statistics [43]. When $\mathcal{Z}(\mathcal{C})$ is non-Abelian, the interpretation is less obvious. In the following, we provide a proof of the fluxon identites from the tube algebra, and comment on the relation with anyon condensation.
@@ C.1.1 Commutative input category
In the commutative case, we can use the mock $S$-matrix $\tilde{s}$ to diagonalize simultaneously all the fusion matrices $N_{i}$. It is a unitary matrix but, unlike the $S$-matrix, it is not symmetric in general (a special case is when the input category is modular, in which case $\tilde{s}$ is a genuine $S$-matrix). Naturally, one would label the matrix elements as $\tilde{s}_{i,j}$ with $i$ and $j\in\mathcal{C}$. However, $\tilde{s}$ is also a unitary transformation between input strings and fluxons (see Eqs. (85) and (86) below). As such, it makes physically more sense to write a matrix element of $\tilde{s}$ as $\tilde{s}_{i,A}$ with rows indexed by input labels $i=1,..,N_{\mathcal{C}}$ and columns by fluxons $A\in\mathcal{F}$ with $N_{\mathcal{F}}=N_{\mathcal{C}}$. The matrix $\tilde{s}$ can be chosen such that its first row only contains strictly positive elements, in which case, $\tilde{s}_{1,A}=\sqrt{d_{A}/\mathcal{D}}=\sqrt{S_{\mathbf{1},A}}$. The mock $S$-matrix is not yet unique as its columns can always be permuted. By convention, we choose that the first column corresponds to the vacuum $A=\mathbf{1}$ of the output category so that $\tilde{s}_{i,\mathbf{1}}=d_{i}/\sqrt{\mathcal{D}}$. Examples of mock $S$-matrix are given in Secs. VI.1, VI.2 and VI.4.
@@
The action of the modular matrices $S$ and $T$ on the tubes is well-known: they act trivially on the empty tube (see Sec. 28.3 in [22]). So that
@@ C.1.2 Noncommutative input category
In the noncommutative case, we also have that the tube algebra restricted to the $11$ sector is the fusion algebra of the input category and that it decouples from the rest of the tube algebra. However, it is noncommutative and there is no mock-$S$ matrix. A consequence is that the number of tubes $N_{\mathcal{C}}$ is strictly larger than the vector space dimension $V_{11}$ which is strictly larger than the number of fluxons $N_{\mathcal{F}}$.
@@
as $(M^{1}_{A,111})_{a,b}=(\Omega^{1}_{A,111})_{a,b}=\delta_{a,b}$ (see above Eq. (40) in Ref. [14] or below Eq. (52) and Eq. (53) in Ref. [12]). Therefore $Q_{1}$ is also the projector onto the $11$ sector, i.e. the empty tube $Q^{1}_{111}$, from which the fluxon identities (92) follow as in the previous subsection.